OpenStudy (anonymous):

What is the least positive integer x for which 12x is the cube of an integer?

OpenStudy (anonymous):

@phi

OpenStudy (kinggeorge):

First, you'll notice that \(12=2^2\cdot3\). So if \(12x=n^3\) for some integer \(n\), then \(3\) must divide \(n^3\), so \(3\) must divide \(n\). Make sense so far?

OpenStudy (phi):

write 12x as 2*2*3*x now you need 3 copies of everything... 3 2's and three 3's what should x be ?

OpenStudy (anonymous):

so 2*2*2*3*3*3*x?

OpenStudy (phi):

you want x so that you end up with 2*2*2*3*3*3

OpenStudy (anonymous):

ok totaly confused. 2*2*2*3*3*3= 216?

OpenStudy (phi):

the idea is if you have 2*2*3 * more_stuff to make that a "cube" you want 3 of each number. you have 2*2 , and if you multiply by 2 you would get 2*2*2* 3 but you need two more 3's, so you should multiply by 3*3 also all together you multiply by some number to change 2*2*3 to 2*2*2*3*3*3 . what number is that ?

OpenStudy (phi):

I meant , what number do we multiply 2*2*3 to get 2*2*2*3*3*3 ? don't you think we should multiply by 2*3*3 ?

OpenStudy (anonymous):

no?

OpenStudy (phi):

Let's start over we have 2*2*3*x we want 2*2*2* 3*3*3 what is x?

OpenStudy (anonymous):

2*3*3?

OpenStudy (phi):

yes, now change that into a number = 18 x is 18

OpenStudy (anonymous):

so the answer would be 18?

OpenStudy (phi):

yes

OpenStudy (anonymous):

THANK YOU!! can u help with one more?